OpenStudy (anonymous):

Solve 6 and one-ninth minus 3 and one-third . Show your work.

OpenStudy (anonymous):

1/3=3/9 6 1/9 = 55/9 3 3/9 = 30/9 55/9 - 30/9 = 25/9 25/9 = 2 7/9